OpenStudy (anonymous):

Use the Pythagorean Theorem to determine the distance between points A(5, 3) and B(-3, -3). 1) 36 2) 10 3) 64 4) 100

OpenStudy (anonymous):

= \[\sqrt{(5-(-3))^2 + (3-(-3))^2} = \sqrt{8^2 + 6^2} = 10\]